what is no of zeros at end of 100!

Grade:12

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear rahul,

100! = 1 x 2 x 3 x ... x 9 x 10 x ... x 15 x 20 x .... 90 x ... x 95 ... x 100

Now each of 10, 20, 30 ie.e every multiple of ten will add one zero, and hundred will add one extra. Hence total zeros = 11 (10 multiples and one extra for 100)
Now each pair of 2,5 ... 12,15... and so on will give one zero. there are 10 such pairs. So Ten more zeroes. But 75 x 72 = 5400 gives two zeroes, hence total zeroes from ths step is also 11.

now we have to account for those products groups which still leave a number 5 in tens place, this is because we have several numbers (10 total) ending in 4 which multiplied with these can give more zeroes. Remember we have already counted the 2's and the 5's in units place.

25 x 22 = 550, which multiplied by any number with 4 in units place will give one more zero
50 , which multiplied by any number with 4 in units place will give one more zero
this means two more zeroes.

Hence total number of zeroes = 11 + 11 + 2 = 24

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

Sudheesh Singanamalla
13 years ago

i think there is a much simpler approach to this problem.

now we get 11 zeroes from numbers such as 10,20,30,...........,90,100)

now a zero can be added if we have a 2 and a 5 i.e, 2 x 5 = 10.

how many such 2s & 5s we get -
between 1 to 10 we get one set of 2 & 5.
similarly we get one set each - from 11 to 20, from 31 to 40, from 51 to 60, from 61 to 70, from 81 to 90 & from 91 to 10.
total 7 such sets or 7 zeroes can be added.

now 21 to 30, 41 to 50 & 71 to 80 are different.
from 21 to 30 we get two 5s from 25. hence we will get two sets of 2 & 5.
from 41 to 50 we get one 5 from 45 & one more 5 from 50. hence we get two sets of 2 & 5.
from 71 to 80 we get two 5s from 75 or 5 x 5 x 3. hence we get two sets of 2 & 5.
total 6 sets or 6 zeroes can be added.

hence total no of zeroes = 11 + 7 + 6 = 24.